. The paired data consist of the temperature (in °F) on randomly chosen days and the amount a Rose plant grew (in millimeters).

Temp

Growth

65

81

83

93

61

76

72

89

50

65

91

87

STATDISK Results

Correlation Results:

Correlation Coeff, r: 0.86615

Critical r:            ±0.81140

P-Value (two-tailed): 0.02568

Regression Results:

Y= b0 + b1x:

Y Intercept, b0:      40.27844

Slope, b1:            0.59083

  1. Use the STATDISK results provided to construct the regression equation y=mx+b, where m is slope and b is the y-intercept . Round all values three decimal places.                           
  1. Use the regression equation to predict the expected growth when the temperature is 75o F. Round the value to one decimal place. Substitute 75 degrees in for x.
  1. Find r2.

  1. What does r2 tell us? (Use a complete sentence!)
Answer #1

Solution-a:

From regresssion equation

y=0.59083*x+40.27844

  1. Use the regression equation to predict the expected growth when the temperature is 75o F. Round the value to one decimal place. Substitute 75 degrees in for x.

y=0.59083*x+40.27844

For c=75 we get

y=0.59083*75+40.27844

y= 84.59069

y=84.6

Predicted ]expected growth=84.6

  1. Find r2.

r^2=r*r

= 0.86615* 0.86615

=0.7502158

  1. What does r2 tell us? (Use a complete sentence!)

R sq=0.7502158

=0.7502158*100

=75.02%

75.02% variaition in growth is explained by temperature

explained avriance by model=75.02%

unexplained variance=100-75.02=24.98%